OpenStudy (anonymous):

does the equation y=5x represent exponential growth or decay?

jimthompson5910 (jim_thompson5910):

hint: in general for equations in the form y = a*b^x, you will a) have exponential decay if b < 1 OR b) have exponential growth if b > 1

jimthompson5910 (jim_thompson5910):

Another hint: \[\large y = 5^x\] is the same as \[\large y = 1*5^x\]

OpenStudy (anonymous):

So it would be decay?

jimthompson5910 (jim_thompson5910):

what are the values of 'a' and b in this case?

OpenStudy (anonymous):

y=5

jimthompson5910 (jim_thompson5910):

I'll start you off a = 1 b = ??

OpenStudy (anonymous):

im not really sure.. ive never done this before. is it y=1*5^x

jimthompson5910 (jim_thompson5910):

that's the more detailed equation compare that to y = a*b^x notice how a = 1 what does b equal?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

so its growth?

jimthompson5910 (jim_thompson5910):

b = 5 since 5 > 1, we know that b > 1 therefore, we have exponential growth

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome